52 inch tv cost 1,995 which was 5% less than the year before, how much was it the year before
Answer by Theo(13342)

it was 1995 last year.
it's 5% less this year.

let x = last year's price.
last year's price minus 5% of last year's price becomes x - .05 * x = 1995
factor out the x to get (1 - .05) * x = 1995
simplify to get .95 * x = 1995
solve for x to get x = 1995 / .95 = 2100.

it was 2100 last year.
5% of 2100 = 105
subtract 105 from 2100 and you get 1995.

your solution is that it was 2100 last year.